# Complex-Valued Holographic Radiance Fields

### v1

@@ 6.4. Qualitative Analysis
Fig. 7 compares simulated defocus blur across three methods: optimization, 3DGS + UNet, and ours. Our method generates perceptually more plausible defocus blur, similar to the optimization-based result without per-view recalculation. 3DGS + UNet‘s result suffers from structured, fringing effect in defocus blur, which is commonly found in learned methods (Shi et al., 2021; Chen et al., 2025).
@@
We evaluate novel view image quality on the NeRF Synthetic, LLFF, and Mip-NeRF 360 datasets—shown in Fig. 8, Fig. 9, and Fig. 10, respectively—using both simulation and experimental captures. For NeRF Synthetic dataset, our method achieves visually consistent results across different datasets and on-par image quality compared with the baseline.
For LLFF dataset, our method preserves image quality effectively in regions with stable depth content, particularly for background elements. However, artifacts become apparent in areas containing near-field objects that exhibit rapid motion across viewpoints, such as foreground furniture and close-range structural elements. These artifacts manifest as phase inconsistencies and visual distortions, primarily in regions where objects undergo significant displacement due to motion parallax.
For Mip-NeRF 360 dataset, the most significant quality loss of our method occurs in distant background regions that experience substantial motion across large viewpoint changes, such as far-field vegetation and architectural structures. Objects at intermediate and near depths maintain reasonable reconstruction quality. This depth-dependent performance indicates that limitations arise from phase discontinuities caused by motion parallax rather than deficiencies in our complex-valued representation, as evidenced by the preserved quality in spatially coherent depth regions. For more simulated and captured result, please refer to Supplementary Sec. D.
@@
To display our complex-valued holograms on a phase-only , we leverage a propagation-based optimization approach that converts amplitude-phase into phase-only representations. Given our complex hologram $P=Ae^{j\varphi}$ with amplitude $A$ and phase $\varphi$, we aim to obtain a corresponding phase-only hologram $P_{\varphi}=e^{j\varphi_{\text{opt}}}$ where the amplitude is unity and the phase $\varphi_{\text{opt}}$ is optimized. We propagate both holograms to the same set of depth planes and minimizes their propagation differences. For each depth plane $\Pi_{l}$, we compute the propagated complex fields
@@
A future version of our method could replace the optimization process with learned approaches, such as U-Net and more, to enable more efficient conversion.
@@ 7.1. Limitations
Speed Trade-offs Between CGH and .
Although our method renders 3D holograms effectively, the rendering process involves splatting Gaussian primitives onto multiple depth planes, followed by wave propagation using the band-limited . The band-limited method is not as efficient as differential splatting, particularly at high resolutions, as it involves multiple high-resolution s to compute accurate wave propagation, which inevitably increases the computation and the memory cost. A more efficient light propagation method is required to further reduce the memory cost and match the rendering speed of intensity-based .
@@
Our method’s reliance on Gaussian primitives for smooth spatial interpolation creates limitations in scenes with significant motion parallax. Objects at dramatically different depths exhibit varying displacements during viewpoint changes, generating sharp phase discontinuities that violate the smoothness assumptions of intensity-based representations. This explains our degraded results on in-the-wild datasets with large depth variations. Future works could explore hierarchical Gaussians (Wang et al., 2025) to adaptively model phase discontinuities at multiple scales.
@@ 7.2. Future Work
A future variant could explore estimating complex-valued holographic radiance fields that dynamically adapt to pupil movements under novel views (Chakravarthula et al., 2022). Moreover, our method inherently contains light field information within the complex-valued 3D representation. With appropriate training objectives, a future extension of our approach could directly support real-time 4D CGH rendering (Kim et al., 2024) without requiring explicit light field extraction.
Time-multiplexed Holography.
By performing multiple rounds of optimization on the same scene, it is possible to generate distinct complex-valued radiance variants. These can be displayed in a time-multiplexed fashion using high-speed  (Choi et al., 2022) to further enhance the image quality and compensate for the phase discontinuities.
@@
Physically Accurate Geometric Alignment.
In the physical world, light amplitude is naturally characterized by its spectral distribution, not through mathematical constructs like . To maintain geometrically aligned phase, our complex-valued holographic radiance field uses only RGB triplets to define the intrinsic amplitude of Gaussians to align with the intrinsic phase values. Despite this simplification, our experiments demonstrate that complex-valued holographic radiance field preserves view-dependent reflectance properties of 3D scenes without requiring high-dimensional view-dependent features. Compared to , which typically contains 48 parameters, our method requires only 6 (3 for amplitude, 3 for phase), resulting in reduced model size and memory footprint. Future work is required to investigate the theoretical foundations of this observation and explore if physically plausible parameterizations can further enhance the expressiveness of the radiance fields.
@@ Conclusion.
Our work introduces a novel approach to support holographic representations in Gaussian Splatting by utilizing our novel 3D complex-valued Gaussians as primitives. We demonstrate that this method leads to a coherent novel view synthesis that is faithful to the geometry of a 3D scene in terms of both phase and amplitude. These findings from our work demonstrate the potential for more physically accurate 3D scene representations, helping holograms to replace 2D images and holographic displays to be compatible and practical as 3D screens in the future.
@@ Appendix A Hardware Image
Fig. 11 shows the photograph of the holographic display prototype we used in this paper. The optical path of our display prototype begins with a laser light source (LASOS MCS4), which integrates three individual laser lines. The emitted light from a single-mode fibre is collimated using a Thorlabs LA1708-A plano-convex lens with a 200 mm focal length. This linearly polarized, collimated beam is then directed by a beamsplitter (Thorlabs BP245B1) toward our phase-only , the Jasper JD7714 (2400×4094, 3.74 $\mu$m). The modulated beam subsequently passes through a lens system comprising Thorlabs LA1908-A and LB1056-A, with focal lengths of 500 mm and 250 mm, respectively. Following this, a pinhole aperture (Thorlabs SM1D12) is positioned at the focal plane of the lenses. Finally, we capture the holographic reconstructions using a lensless image sensor (Point Grey GS3-U3-23S6M-C USB 3.0), which is mounted on an X-stage (Thorlabs PT1/M) with a travel range of 0 to 25 mm and a positioning precision of 0.01 mm.
@@ B.1. Wave Propagation From a Single 3D Gaussian Primitive
In a pure intensity world, to account for fore-ground occlusion, the frontmost Gaussian along the line of sight tends to occlude anything behind it, and people apply alpha-blending to represent transparent pixels. For simulating physical wave propagation in a hologram, occlusion is subtler—each ”surface” can block or partially block wavefronts behind it. One approach to accurately model occlusion is to reconstruct a triangle surface mesh by connecting the neighboring points of Gaussians and performing ray casting from each point to the hologram plane $P$. The wavefronts carried by rays that hit the surface mesh are excluded from the hologram calculation. A binary visibility mask can be generated from this process to mimic foreground-occlusion at the hologram plane. However, both point-based wave simulation and ray-triangle detection of the entire scene are computationally expensive, making them a less favorable choice in real-time inference. As illustrated in Fig. 12, in our implementation, we use alpha-blending with transmittance to efficiently approximate occlusion effects, each 3D Gaussian $\mathcal{G}_{n}$ can be considered as a small volume in $\Omega$ that emits a coherent wavefield. rather than the more computationally expensive ray-casting approach described above.
@@
Appendix C Differentiable Complex-Valued Gaussian Rendering
@@
C.1. Notation
  • $\mathbf{x}_{n}\in\mathbb{R}^{3}$ 3D Gaussian center positions in world space
  • $\mathbf{x}_{n}^{\text{proj}}\in\mathbb{R}^{2}$ Projected 2D Gaussian centers in screen space
  • $\mu_{x},\mu_{y}$ Components of the 2D mean position
  • $\Sigma\in\mathbb{R}^{3\times 3}$ 3D covariance matrix in world space
  • $\Sigma^{\prime}\in\mathbb{R}^{2\times 2}$ 2D covariance matrix in screen space
  • $\Sigma^{\prime-1}_{ij}$ Elements of the inverse 2D covariance matrix (where $i,j\in\{0,1\}$)
  • $d_{x}=x-\mu_{x},d_{y}=y-\mu_{y}$ Distance from pixel to Gaussian center
  • $\mathbf{p}_{n,l}$ Plane assignment probability of Gaussian $n$ for plane $l$ (controls contribution to each rendering plane)
  • $\boldsymbol{\varphi}_{n}$ Phase value for Gaussian $n$ (controls complex field phase)
  • $\boldsymbol{\alpha}_{n}$ Opacity value for Gaussian $n$
  • $\boldsymbol{\alpha}_{n}^{\text{base}}$ Base opacity value before applying the Gaussian function
  • $T_{n}$ Accumulated transmittance at Gaussian $n$ (product of previous transparency values)
  • $z_{\text{norm}}$ Normalized depth value between 0 and 1
C.2. Optimization and Densification Algorithm
Our optimization and densification algorithms extend the standard  (Kerbl et al., 2023) approach to handle complex-valued parameters and multi-plane assignments. The key modifications include phase parameter optimization, plane probability learning with , and complex-valued loss computation as summarized in Algorithm 1.
@@ C.2. Optimization and Densification Algorithm
3:$M\leftarrow$ SfM Points $\triangleright$ Positions
4:$\mathbf{S},\mathbf{C},\boldsymbol{\alpha}\leftarrow$ InitAttributes() $\triangleright$ Scales, Amplitudes, Opacities
5:$\boldsymbol{\varphi}\leftarrow$ InitPhase() $\triangleright$ Phase Parameters
6:$\mathbf{p^{\prime}}\leftarrow$ InitPlaneProbs($L$) $\triangleright$ Plane Assignment Logits
7:$i\leftarrow 0$triangleright Iteration Count
@@
10:    $\mathbf{p}\leftarrow$ STE($\mathbf{p^{\prime}}$) $\triangleright$ Hard Assignment via STE
11:    $U_{\text{complex}}\leftarrow$ ComplexRasterize($M,\mathbf{S},\mathbf{C},\boldsymbol{\alpha},\boldsymbol{\varphi},\mathbf{p},V$) $\triangleright$ Alg. 2
@@
15:    $M,\mathbf{S},\mathbf{C},\boldsymbol{\alpha},\boldsymbol{\varphi},\mathbf{p^{\prime}}\leftarrow$ $\nabla\mathcal{L}$) $\triangleright$ Backprop & Step
@@
17:        for all Gaussians $(\mu,\Sigma,c,\alpha,\varphi,p^{\prime})$ in $(M,\mathbf{S},\mathbf{C},\boldsymbol{\alpha},\boldsymbol{\varphi},\mathbf{p^{\prime}})$ do
18:           if $\alpha<\epsilon$ or IsTooLarge($\mu,\Sigma,c,\alpha$then $\triangleright$ Pruning
@@
23:                  SplitGaussian($\mu,\Sigma,c,\alpha,\varphi,p^{\prime}$)
24:               $\triangleright$ Under-reconstruction
25:                  CloneGaussian($\mu,\Sigma,c,\alpha,\varphi,p^{\prime}$)
@@
The key extensions to the standard algorithm include: (1) Phase parameter $\boldsymbol{\varphi}$ initialization with random values in $[0,2\pi)$; (2) Plane assignment probabilities $\mathbf{p^{\prime}}$ initialized uniformly and converted to hard assignments via during forward pass; (3) Complex-valued rasterization producing multi-plane complex fields; (4) Hologram generation through multi-plane propagation; (5) Phase-aware densification where both split and clone operations preserve phase relationships by adding small random perturbations to maintain coherent interference patterns.
C.3. Details of the Rasterizer
@@
2:$M,\mathbf{S}$: Gaussian means and covariances in world space
3:$\mathbf{C},\boldsymbol{\alpha},\boldsymbol{\varphi},\mathbf{p}$: Amplitudes, opacities, phases, and plane assignments
@@
6:function ComplexRasterize($w,h,M,\mathbf{S},\mathbf{C},\boldsymbol{\alpha},\boldsymbol{\varphi},\mathbf{p},V$)
@@
16:           for all Pixels $pix$ in $t$ do
17:               $T_{\text{acc}}\leftarrow 1.0$triangleright Transmittance
18:               $\text{real}_{\text{acc}},\text{imag}_{\text{acc}}\leftarrow 0.0,0.0$triangleright Complex Accumulation
@@
21:                  $w_{\text{plane}}\leftarrow\mathbf{p}_{g,l}$triangleright Plane Assignment
@@
25:                  $\alpha_{\text{eff}}\leftarrow\boldsymbol{\alpha}_{g}\cdot\exp(-\text{power}_{g,\text{pix}})\cdot w_{\text{plane}}$triangleright Effective Alpha
26:                  $\text{scale}\leftarrow\mathbf{C}_{g}\cdot\alpha_{\text{eff}}\cdot T_{\text{acc}}$triangleright Complex Scale
27:                  $\text{real}_{\text{acc}}\leftarrow\text{real}_{\text{acc}}+\text{scale}\cdot\cos(\boldsymbol{\varphi}_{g})$triangleright Real Component
28:                  $\text{imag}_{\text{acc}}\leftarrow\text{imag}_{\text{acc}}+\text{scale}\cdot\sin(\boldsymbol{\varphi}_{g})$triangleright Imaginary Component
29:                  $T_{\text{acc}}\leftarrow T_{\text{acc}}\cdot(1-\alpha_{\text{eff}})$triangleright Update Transmittance

### v2

@@ 6.4. Qualitative Analysis
Fig. 7 compares simulated defocus blur across three CGH methods: optimization, 3DGS + UNet, and ours. Our method generates perceptually more plausible defocus blur, similar to the optimization-based result without per-view recalculation. 3DGS + UNet’s result suffers from a structured, fringing effect in defocus blur, which is commonly found in learned CGH methods (Shi et al., 2021; Chen et al., 2025).
@@
We evaluate novel view image quality on the NeRF Synthetic, LLFF, and Mip-NeRF 360 datasets—shown in Fig. 8, Fig. 9, and Fig. 10, respectively—using both simulation and experimental captures. For the NeRF Synthetic dataset, our method achieves visually consistent results across different datasets and on-par image quality compared to the baseline.
For the LLFF dataset, our method preserves image quality effectively in regions with stable depth content, particularly for background elements. However, artifacts become apparent in areas containing near-field objects that exhibit rapid motion across viewpoints, such as foreground furniture and close-range structural elements. These artifacts manifest as phase inconsistencies and visual distortions, primarily in regions where objects undergo significant displacement due to motion parallax.
For the Mip-NeRF 360 dataset, the most significant quality loss of our method occurs in distant background regions that experience substantial motion across large viewpoint changes, such as far-field vegetation and architectural structures. Objects at intermediate and near depths maintain reasonable reconstruction quality. This depth-dependent performance indicates that limitations arise from phase discontinuities caused by motion parallax rather than deficiencies in our complex-valued representations, as evidenced by the preserved quality in spatially coherent depth regions. For more simulated and captured results, please refer to Supplementary Sec. G.
@@
To display our complex-valued holograms on a phase-only SLM, we leverage a propagation-based optimization approach that converts amplitude-phase into phase-only representations. Given our complex hologram $P=Ae^{j\varphi}$ with amplitude $A$ and phase $\varphi$, we aim to obtain a corresponding phase-only hologram $P_{\varphi}=e^{j\varphi_{\text{opt}}}$ where the amplitude is unity and the phase $\varphi_{\text{opt}}$ is optimized. We propagate both holograms to the same set of depth planes and minimize their propagation differences. For each depth plane $\Pi_{l}$, we compute the propagated complex fields
@@
Future variants could adopt learned approaches, such as U-Net and others, to replace optimization and enable more efficient conversion.
@@ 7.1. Limitations
Wave Propagation Overhead.
Although our method renders 3D holograms effectively, the rendering process involves splatting Gaussian primitives onto multiple depth planes, followed by wave propagation using the band-limited ASM. The band-limited ASM method is not as efficient as differential splatting, particularly at high resolutions, as it involves multiple high-resolution FFTs to compute accurate wave propagation, which inevitably increases the computation and the memory cost. A more efficient light propagation method (Zhan et al., 2025a) is required to further reduce the memory cost and match the rendering speed of intensity-based 3DGS.
@@
Our method’s reliance on Gaussian primitives for smooth spatial interpolation creates limitations in scenes with significant motion parallax. Objects at dramatically different depths exhibit varying displacements during viewpoint changes, generating sharp phase discontinuities that violate the smoothness assumptions of intensity-based representations. This explains our degraded results on in-the-wild datasets with large depth variations. Future work could explore hierarchical Gaussians (Wang et al., 2025) to adaptively model phase discontinuities at multiple scales.
@@
Learned Plane Assignment.
The plane assignment probabilities in our method are a computational latent variable rather than a physically-accurate depth labeling. The consistency across interpolated novel viewpoints is learned implicitly from the training view distribution, rather than guaranteed by a physical model. While effective in practice, this approach is a pragmatic compromise. Future work should explore more elegant and physically grounded alternatives.
@@ 7.2. Future Work
A future variant could explore estimating complex-valued holographic radiance fields that dynamically adapt to pupil movements under novel views (Chakravarthula et al., 2022). Moreover, our method inherently contains light field information within the complex-valued 3D representations. With appropriate training objectives, a future extension of our approach could directly support real-time 4D CGH rendering (Kim et al., 2024) without requiring explicit light field extraction.
@@
Physically-informed Color Representation.
In the physical world, light amplitude is naturally characterized by its spectral distribution, not through mathematical constructs like Spherical Harmonic (SH). To maintain scene geometry-aware representations, our complex-valued holographic radiance field uses only RGB triplets to define the intrinsic amplitude of Gaussians to align with the intrinsic phase values. Despite this simplification, our experiments demonstrate that complex-valued holographic radiance field preserves view-dependent reflectance properties of 3D scenes without requiring high-dimensional view-dependent features. Compared to SH, which typically contains 48 parameters, our method requires only 6 (3 for amplitude, 3 for phase), resulting in reduced model size and memory footprint. Future work could examine the theoretical basis of this observation and assess whether physically plausible parameterizations further enhance radiance field expressiveness.
@@ Conclusion.
Our work introduces a novel approach to support holographic representations in Gaussian Splatting by utilizing our novel 3D complex-valued Gaussians as primitives. We demonstrate that this method leads to a coherent novel view synthesis that is faithful to the geometry of a 3D scene in terms of both amplitude and phase. These findings from our work demonstrate the potential for more physically-accurate 3D scene representations, helping holograms to replace 2D images and holographic displays to be compatible and practical as 3D screens in the future.
Acknowledgements.
The authors thank Dr. Josef Spjut and Dr. Mike Roberts for providing valuable suggestions in the early phases. Seung-Hwan acknowledges funding from the National Research Foundation of Korea (NRF) grants funded by the Korea government (MSIT) (RS-2024-00438532, RS-2023-00211658) and the Ministry of Education through the Basic Science Research Program (2022R1A6A1A03052954), as well as grants from the Institute of Information & Communications Technology Planning & Evaluation (IITP) funded by the Korea government (MSIT) (No. RS-2024-0045788) and the IITP-ITRC (Information Technology Research Center) program (IITP-2026-RS-2024-00437866).
@@ Appendix A Hardware Image
[…]
@@ B.1. Wave Propagation From a Single 3D Gaussian Primitive
In a pure intensity world, to account for fore-ground occlusion, the frontmost Gaussian along the line of sight tends to occlude anything behind it, and people apply alpha-blending to represent transparent pixels. For simulating physical wave propagation in a hologram, occlusion is subtler—each ”surface” can block or partially block wavefronts behind it. One approach to accurately model occlusion is to reconstruct a triangle surface mesh by connecting the neighboring points of Gaussians and performing ray casting from each point to the hologram plane $P$. The wavefronts carried by rays that hit the surface mesh are excluded from the hologram calculation. A binary visibility mask can be generated from this process to mimic foreground-occlusion at the hologram plane. However, both point-based wave simulation and ray-triangle detection of the entire scene are computationally expensive, making them a less favorable choice in real-time inference. As illustrated in Fig. 12, in our implementation, we use alpha-blending with transmittance to efficiently approximate occlusion effects, each 3D Gaussian $\mathcal{G}_{n}$ can be considered as a small volume in $\Omega$ that emits a coherent wavefield, rather than the more computationally expensive ray-casting approach described above.
@@
Appendix C Training Camera Viewpoint Sampling
During training, camera viewpoints are sampled differently depending on the dataset characteristics to ensure the learned scene representation supports consistent rendering across viewpoints within each dataset’s coverage.
For the NeRF Synthetic dataset, we randomly sample camera poses facing towards the object within the Blender coordinate system, maintaining the same radial distance distribution as the original dataset while randomizing the azimuth and elevation angles. This generates a continuous distribution of training viewpoints around the object.
For the LLFF dataset, we directly use the existing camera viewpoints provided in the dataset, as the captured views already provide sufficient coverage of the indoor scenes with manageable motion parallax.
For the Mip-NeRF 360 dataset, we employ a two-stage approach due to complex field discontinuities challenges arising from extreme motion parallax in the original wide-baseline captures. First, we train an intensity-based 3DGS scene using the provided dataset images. We uniformly sample 480 camera views covering a full 360° rotation, corresponding to 0.75° angular spacing for training, and evaluate novel-view rendering on the intermediate unseen viewpoints between sampled angles. We then use this trained 3DGS scene as a reference to randomly sample camera viewpoints facing towards the scene center, similar to the NeRF Synthetic sampling strategy. As a way round method, this approach generates training views with reduced inter-view motion parallax in the near scene compared to the original Mip-NeRF 360 captures, mitigating the severe complex field discontinuities that would otherwise cause substantial image quality degradation, but it does not help with the far field.
Appendix D Differentiable Complex-Valued Gaussian Rendering
@@
D.1. Notation
  • $\mathbf{x}_{n}\in\mathbb{R}^{3}$ - 3D Gaussian center positions in world space
  • $\mathbf{x}_{n}^{\text{proj}}\in\mathbb{R}^{2}$ - Projected 2D Gaussian centers in screen space
  • $\mu_{x},\mu_{y}$ - Components of the 2D mean position
  • $\Sigma\in\mathbb{R}^{3\times 3}$ - 3D covariance matrix in world space
  • $\Sigma^{\prime}\in\mathbb{R}^{2\times 2}$ - 2D covariance matrix in screen space
  • $\Sigma^{\prime-1}_{ij}$ - Elements of the inverse 2D covariance matrix (where $i,j\in\{0,1\}$)
  • $d_{x}=x-\mu_{x},d_{y}=y-\mu_{y}$ - Distance from pixel to Gaussian center
  • $\boldsymbol{\rho}_{n,l}$ - Plane assignment probability of Gaussian $n$ for plane $l$ (controls contribution to each rendering plane)
  • $\boldsymbol{\varphi}_{n}$ - Phase value for Gaussian $n$ (controls complex field phase)
  • $\boldsymbol{\alpha}_{n}$ - Opacity value for Gaussian $n$
  • $\boldsymbol{\alpha}_{n}^{\text{base}}$ - Base opacity value before applying the Gaussian function
  • $T_{n}$ - Accumulated transmittance at Gaussian $n$ (product of previous transparency values)
  • $z_{\text{norm}}$ - Normalized depth value between 0 and 1
D.2. Optimization and Densification Algorithm
Our optimization and densification algorithms extend the standard 3DGS (Kerbl et al., 2023) approach to handle complex-valued parameters and multi-plane assignments. The key modifications include phase parameter optimization, plane probability learning with STE, and complex-valued loss computation as summarized in Algorithm 1.
@@ D.2. Optimization and Densification Algorithm
[…]
@@
10:  $\boldsymbol{\rho}\leftarrow$ STE($\boldsymbol{\rho}^{\prime}$) $\triangleright$ Hard Assignment via STE
11:  $U_{\text{complex}}\leftarrow$ ComplexRasterize($M,\mathbf{S},\mathbf{C},\boldsymbol{\alpha},\boldsymbol{\varphi},\boldsymbol{\rho},V$) $\triangleright$ Alg. 2
@@
15:  $M,\mathbf{S},\mathbf{C},\boldsymbol{\alpha},\boldsymbol{\varphi},\boldsymbol{\rho}^{\prime}\leftarrow$ Adan($\nabla\mathcal{L}$) $\triangleright$ Backprop & Step
@@
17:   for all Gaussians $(\mu,\Sigma,c,\alpha,\varphi,\rho^{\prime})$ in $(M,\mathbf{S},\mathbf{C},\boldsymbol{\alpha},\boldsymbol{\varphi},\boldsymbol{\rho}^{\prime})$ do
18:     if $\alpha<\epsilon$ or IsTooLarge($\mu,\Sigma,c,\alpha$) then $\triangleright$ Pruning
@@
23:       SplitGaussian($\mu,\Sigma,c,\alpha,\varphi,\rho^{\prime}$)
[…]
25:       CloneGaussian($\mu,\Sigma,c,\alpha,\varphi,\rho^{\prime}$)
@@
The key extensions to the standard algorithm include: (1) Phase parameter $\boldsymbol{\varphi}$ initialization with random values in $0,2\pi)$; (2) Plane assignment probabilities $\boldsymbol{\rho}^{\prime}$ initialized uniformly and converted to hard assignments via STE during forward pass; (3) Complex-valued rasterization producing multi-plane complex fields; (4) Hologram generation through multi-plane ASM propagation; (5) Phase-aware densification where both split and clone operations preserve phase relationships by adding small random perturbations to maintain coherent interference patterns.
D.3. Details of the Rasterizer
@@
[…]
@@
6:function ComplexRasterize($w,h,M,\mathbf{S},\mathbf{C},\boldsymbol{\alpha},\boldsymbol{\varphi},\boldsymbol{\rho},V$)
@@
16:     for all Pixels pixpix in $t$ do
17:      $T_{\text{acc}}\leftarrow 1.0$ $\triangleright$ Transmittance
18:      $\text{real}_{\text{acc}},\text{imag}_{\text{acc}}\leftarrow 0.0,0.0$ $\triangleright$ Complex Accumulation
@@
21:       $w_{\text{plane}}\leftarrow\boldsymbol{\rho}_{g,l}$ $\triangleright$ Plane Assignment
@@
25:       $\alpha_{\text{eff}}\leftarrow\boldsymbol{\alpha}_{g}\cdot\exp(-\text{power}_{g,\text{pix}})\cdot w_{\text{plane}}$ $\triangleright$ Effective Alpha
26:       $\text{scale}\leftarrow\mathbf{C}_{g}\cdot\alpha_{\text{eff}}\cdot T_{\text{acc}}$ $\triangleright$ Complex Scale
27:       $\text{real}_{\text{acc}}\leftarrow\text{real}_{\text{acc}}+\text{scale}\cdot\cos(\boldsymbol{\varphi}_{g})$ $\triangleright$ Real Component
28:       $\text{imag}_{\text{acc}}\leftarrow\text{imag}_{\text{acc}}+\text{scale}\cdot\sin(\boldsymbol{\varphi}_{g})$ $\triangleright$ Imaginary Component
29:       $T_{\text{acc}}\leftarrow T_{\text{acc}}\cdot(1-\alpha_{\text{eff}})$ $\triangleright$ Update Transmittance